# Composition Theorems for Multiple Differential Privacy Constraints

### 21 Mar 2026 v1

@@ I-A Differential Privacy and f-DP
Def. 1 could be alternatively formulated as a constraint on the trade-off function of the hypothesis test employed by the adversary [25, 17]. Motivated by this, [8] introduced $f$-DP. In $f$-DP, trade-off functions $f$ have infinite degrees of freedom and can precisely describe the privacy of a mechanism. This is in contrast with $(\varepsilon,\delta)$-DP where the privacy level is expressed with two parameters. One interesting feature of the $f$-DP framework is the equivalence of a symmetric trade-off function $f$ and a (potentially infinite) collection of $(\varepsilon,\delta)$-DP guarantees. This equivalence suggests that an $f$-DP guarantee can be approximated by multiple DP constraints and motivates the study of such settings in this paper. By using multiple DP constraints, we can approach an arbitrary $f$-DP constraint and approximate its $k$-fold composition.
@@ I-B Composition Theorems
In most differentially private applications, the same database is reused for multiple queries as opposed to requesting just a single query. Privacy degrades with composition and it is crucial to quantify the loss closely. The composition theorems characterize this privacy loss. Dwork et al. [10] provide an initial bound on $k$ compositions of $(\varepsilon,\delta)$-DP and show that it is $(k\varepsilon,k\delta)$-DP. Since then, there has been a large body of work on composition theorems for differential privacy [10, 11, 12, 23, 1, 7, 17, 8, 19, 2, 14, 18, 27, 13, 15]. Some recent works have taken advantage of the hypothesis testing point of view of DP [25, 17, 23, 13]. Kairouz et al. [17] compute the exact $k$-fold composition region for a single $(\varepsilon,\delta)$-DP constraint in the adaptive setting. Ghazi et al. [13] extend this result by taking total variation privacy into account. This is equivalent to having simultaneous $(\varepsilon,\delta)$-DP and $(0,\eta)$-DP constraints on all $k$ mechanisms. This approach achieves tighter composition results compared to [17] whenever the mechanisms attain a non-trivial total variation constraint. Murtagh et al. [23] prove a tight heterogeneous composition theorem in a similar manner. Dong et al. [8] take a different approach and show a central-limit like composition result for sufficiently smooth trade-off functions, yielding an approximation method for composition.
@@ I-C Our Contributions
We build on the existing composition results and investigate the composition of double-DP constraint mechanisms, which satisfy $(\varepsilon_{1},\delta_{1})$-DP and $(\varepsilon_{2},\delta_{2})$-DP, simultaneously. In Theorems 2 and 3, we establish the exact composition region in this setting. Our result in Theorem 2 reveals the intimate relationship between the composition of double-DP constraint mechanisms and the heterogeneous composition. Thus, we find the exact heterogeneous composition of $n$ $\varepsilon_{1}$-DP mechanisms and $m$ $\varepsilon_{2}$-DP mechanisms in Theorem 1. This result achieves a lower computational complexity compared to [23]. In addition to facilitating Theorem 2, Theorem 1 is of independent interest. Its setting reflects a realistic scenario in which the privacy constraints need to be adjusted at some point to satisfy updated requirements. Finally, in Propositions 1 and 2, we develop a method to approximate $f$-DP composition via double-DP constraint composition, motivated by [8].
@@ II-A Hypothesis Testing and $f$-DP
This interpretation inspired a very general variant of DP, called $f$-DP [8]. Informally, a mechanism $M$ satisfies $f=f(P,Q)$-DP if differentiating two neighboring databases $D_{0}$ and $D_{1}$ through $M(D_{0})$ and $M(D_{1})$ is at least as difficult as differentiating $P$ and $Q$, in a strong sense. Equation (4) indicates that $f$-DP reduces to $(\varepsilon,\delta)$-DP with the appropriate trade of function, i.e. $f=f_{\varepsilon,\delta}$.
@@ II-B Privacy Regions
Privacy regions offer an alternative way to express trade-off functions. They are regions including all achievable error pairs and have one to one correspondence to the trade-off functions.
@@
When we talk about multiple $(\varepsilon_{i},\delta_{i})$-DP constraints, we use the vector notation, and denote it with bold letters. For example, we let $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{n})\in(\mathbb{R}^{+})^{n}$ and $\boldsymbol{\delta}=(\delta_{1},\ldots,\delta_{n})\in[0,1]^{n}$. We denote the privacy region of the intersection of the $n$ $(\varepsilon_{i},\delta_{i})$-DP guarantees by
@@
This intersection region signifies the privacy region of a mechanism that satisfy all $(\varepsilon_{i},\delta_{i})$-DP constraints simultaneously. We study the $k$-fold composition of such mechanisms.
@@ II-C Mixture of Hypothesis Tests
Consider that $f_{m}=\inf_{\hat{\mathcal{H}}_{m}}\left\{\beta_{\mathrm{II}}^{m}|\beta_{\mathrm{I}}^{m}\leq t\right\}$. For $\nu\in\left\{\mathrm{I},\mathrm{II}\right\}$, $\beta_{\nu}(\hat{\mathcal{H}}_{m})=\sum_{i=1}^{n}\alpha_{i}\beta_{\nu}^{i}(\hat{\mathcal{H}}_{m}(\cdot,i)):=\sum_{i=1}^{n}\alpha_{i}\beta_{\nu}^{i}(\hat{\mathcal{H}}^{i})$ by routine arguments. The reformulated constraint becomes $\beta_{\mathrm{I}}^{m}=\sum_{i=1}^{n}\alpha_{i}t_{i}\leq t$, $t_{i}=\beta_{\mathrm{I}}^{i}(\hat{\mathcal{H}}^{i})$. Then observe that $f_{i}(t_{i})=f_{i}(\beta_{\mathrm{I}}^{i}(\hat{\mathcal{H}}^{i}))\leq\beta_{\mathrm{II}}^{i}(\hat{\mathcal{H}}^{i})$, leading to the rewriting of $\beta_{\mathrm{II}}^{m}$ as $\sum_{i=1}^{n}\alpha_{i}f_{i}(t_{i})$. The precise arguments can be found in Appendix -A. ∎
@@
Equation 6 is known in the convex optimization literature as a weighted infimal convolution. Its properties are studied in [4], from which the following corollary becomes apparent.
@@
See Fig. 1 for an example of the mixture of two equally weighted DP regions.
@@ III-A Heterogeneous Composition
We state the theorem on heterogeneous composition $\mathcal{C}_{x,y}\left(\varepsilon_{1},\varepsilon_{2}\right)$ (cf. Def. 4) and the accompanying Algorithm 1.
@@ III-B Composition Theorem for Double-DP Constraints
A trivial upper bound of the double-DP $k$-composition privacy region is the intersection of the privacy regions for each DP guarantee, i.e. $\mathcal{R}_{k}\left(\boldsymbol{\varepsilon},\boldsymbol{\delta}\right)\subseteq\mathcal{R}_{k}\left(\varepsilon_{1},\delta_{1}\right)\cap\mathcal{R}_{k}\left(\varepsilon_{2},\delta_{2}\right)$, known thanks to [17].
@@
Another bound can be derived from [13], leveraging the total variation bound induced by DP. Let $\eta=\delta_{2}+(1-\delta_{2})\frac{e^{\varepsilon}_{2}-1}{e^{\varepsilon}_{2}+1}$. Then, $\mathcal{R}_{k}\left(\boldsymbol{\varepsilon},\boldsymbol{\delta}\right)\subseteq\mathcal{R}_{k}\left((\varepsilon_{1},0),(\delta_{1},\eta)\right)\cap\mathcal{R}_{k}\left(\varepsilon_{2},\delta_{2}\right)$.
Remarks 1 and 2 are baselines for our main composition result.
@@
$P_{i}(x)=\begin{cases}(1-\alpha)P_{i}^{1}(x)&\text{ if }x\in\left\{0,3\right\},\\ \alpha P_{i}^{2}(x)&\text{ if }x\in\left\{1,2\right\},\end{cases}$
@@
See Fig. 2 for comparison of the bounds given by Remarks 1 and 2 with the exact privacy region computed in Theorems 2 and 3.
@@ IV A Method for Approximating $f$-DP Composition
In the following, we algorithmically approximate the $k$-composition of a trade-off function $f$, yielding both lower and upper privacy bounds***The corresponding code is public on github at https://go.epfl.ch/multiDP.. This is done by approximating $f$ itself from below by a double-DP trade-off function $f_{\boldsymbol{\varepsilon}^{-},\boldsymbol{\delta}^{-}}$, and same from above by $f_{\boldsymbol{\varepsilon}^{+},\boldsymbol{\delta}^{+}}$, then computing the $k$-composition of each using Theorems 2 or 3. $f_{\boldsymbol{\varepsilon}^{-},\boldsymbol{\delta}^{-}}$ is defined by
@@ IV-A $f$-DP approximation from below
Below, we define the normal rotation of $f$. It is the rotation of the graph of $f$ by $\frac{\pi}{4}$, changing the axes of the coordinate system from $y=0$ and $x=0$ to y=xy=-x and $y=x$.
@@ IV-C Example: Gaussian Mechanism
In Fig. 3, we obtain empirically close lower and upper bounds on $1$-GDP compositions through double-DP, which will become arbitrarily tight once we generalize our results to $n$-DP composition, as $n$ increases. Note that FFT based methods [19, 14] can already achieve arbitrarily close bounds, but are harder to use and more computationally expensive.
@@ V Concluding Remarks
Building on the framework introduced in Theorem 2, the generalization to $n$ DP constraints seems natural and will be the subject of future work. With $\boldsymbol{j}=(j_{1},\ldots,j_{n})\in(\mathbb{N}^{*})^{n}$, we refer to the privacy region of the heterogeneous composition of $j_{i}$ $\varepsilon_{i}$-DP for $i\in\llbracket 1\mkern 1.5mu,\mkern 1.5mun\rrbracket$ by $\mathcal{C}_{\boldsymbol{j}}\left(\boldsymbol{\varepsilon}\right)$.
Conjecture 1.
Let $\boldsymbol{\varepsilon}=(\varepsilon_{i})_{i=1}^{n}\in(\mathbb{R}^{*}_{+})^{n}$ and $\boldsymbol{\delta}=(\delta_{i})_{i=1}^{n}\in[0,1]^{n}$ such that all $(\varepsilon_{i},\delta_{i})$-DP constraints are active. Then the privacy region of the composition of $k$ mechanisms which are $(\varepsilon_{i},\delta_{i})$-DP for $i\in\llbracket 1\mkern 1.5mu,\mkern 1.5mun\rrbracket$ is
$\mathcal{R}_{k}\left(\boldsymbol{\varepsilon},\boldsymbol{\delta}\right)=\tilde{\delta}\mathcal{R}\left(0,1\right)+(1-\tilde{\delta})\sum_{\begin{subarray}{c}\boldsymbol{j}\in\llbracket 0\mkern 1.5mu,\mkern 1.5muk\rrbracket^{n}\\ \sum_{i=1}^{n}j_{i}=k\end{subarray}}\omega_{\boldsymbol{j}}\mathcal{C}_{\boldsymbol{j}}\left(\boldsymbol{\varepsilon}\right)$
for $\omega_{\boldsymbol{j}}=\binom{k}{j_{1},\ldots,j_{n}}\prod_{i=1}^{n}\alpha_{i}^{j_{i}}$ for some functions $\alpha_{i}$, $\tilde{\delta}$ of $\boldsymbol{\varepsilon}$, $\boldsymbol{\delta}$.

### 15 Apr 2026 v2

@@ I-A Differential Privacy and f-DP
Definition 1 could be alternatively formulated as a constraint on the trade-off function of the hypothesis test employed by the adversary [25, 17]. Motivated by this, [8] introduced $f$-DP. In $f$-DP, trade-off functions $f$ have infinite degrees of freedom and can precisely describe the privacy of a mechanism. This is in contrast with $(\varepsilon,\delta)$-DP where the privacy level is expressed with two parameters. One interesting feature of the $f$-DP framework is the equivalence of a symmetric trade-off function $f$ and a (potentially infinite) collection of $(\varepsilon,\delta)$-DP guarantees. This equivalence suggests that an $f$-DP guarantee can be approximated by multiple DP constraints and motivates the study of such settings in this paper. By using multiple DP constraints, we can approach an arbitrary $f$-DP constraint and approximate its $k$-fold composition.
@@ I-B Composition Theorems
In most differentially private applications, the same database is reused for multiple queries as opposed to just a single query. Privacy degrades with composition and it is crucial to quantify the loss closely. The composition theorems characterize this privacy loss. Dwork et al. [10] provide an initial bound on $k$ compositions of $(\varepsilon,\delta)$-DP and show that it is $(k\varepsilon,k\delta)$-DP. Since then, there has been a large body of work on composition theorems for differential privacy [10, 11, 12, 23, 1, 7, 17, 8, 19, 2, 14, 18, 27, 13, 15]. Some recent works have taken advantage of the hypothesis testing point of view of DP [25, 17, 23, 13]. Kairouz et al. [17] compute the exact $k$-fold composition region for a single $(\varepsilon,\delta)$-DP constraint in the adaptive setting. Ghazi et al. [13] extend this result by taking total variation privacy into account. This is equivalent to having simultaneous $(\varepsilon,\delta)$-DP and $(0,\eta)$-DP constraints on all $k$ mechanisms. This approach achieves tighter composition results compared to [17] whenever the mechanisms attain a non-trivial total variation constraint. Murtagh et al. [23] prove a tight heterogeneous composition theorem in a similar manner. Dong et al. [8] take a different approach and show a central-limit like composition result for sufficiently smooth trade-off functions, yielding an approximation method for composition.
@@ I-C Our Contributions
We build on the existing composition results and investigate the composition of double-DP constraint mechanisms, which satisfy $(\varepsilon_{1},\delta_{1})$-DP and $(\varepsilon_{2},\delta_{2})$-DP, simultaneously. In Theorems 2 and 3, we establish the exact composition region in this setting. Our result in Theorem 2 reveals the intimate relationship between the composition of double-DP constraint mechanisms and the heterogeneous composition. Thus, we find the exact heterogeneous composition of $n$ $\varepsilon_{1}$-DP mechanisms and $m$ $\varepsilon_{2}$-DP mechanisms in Theorem 1. This result achieves a lower computational complexity compared to [23]. In addition to facilitating Theorem 2, Theorem 1 is of independent interest. Its setting reflects a realistic scenario in which the privacy constraints need to be adjusted at some point to satisfy updated requirements such as the number of queries or the privacy budget. Finally, in Propositions 1 and 2, we develop a method to approximate $f$-DP composition via double-DP constraint composition, motivated by [8].
@@ II-A Hypothesis Testing and $f$-DP
This interpretation inspired a very general variant of DP, called $f$-DP [8]. Informally, a mechanism $M$ satisfies $f=f(P,Q)$-DP if differentiating two neighboring databases $D_{0}$ and $D_{1}$ through $M(D_{0})$ and $M(D_{1})$ is at least as difficult as differentiating $P$ and $Q$, in a strong sense. Equation (4) indicates that $f$-DP reduces to $(\varepsilon,\delta)$-DP with the appropriate trade of function, i.e. $f=f_{\varepsilon,\delta}$. Likewise, multiple $(\varepsilon_{i},\delta_{i})$-DP constraints could also be reflected in the $f$-DP framework: let $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{n})\in(\mathbb{R}^{+})^{n}$, $\boldsymbol{\delta}=(\delta_{1},\ldots,\delta_{n})\in[0,1]^{n}$, and
$f_{\boldsymbol{\varepsilon},\boldsymbol{\delta}}(t)=\max_{(\varepsilon,\delta)\in(\boldsymbol{\varepsilon},\boldsymbol{\delta})}f_{\varepsilon,\delta}(t)\ \forall t\in[0,1].$ (5)
@@ II-B Privacy Regions
Privacy regions offer an alternative way to express trade-off functions. They include all achievable error pairs and have one to one correspondence to the trade-off functions.
@@
We denote the privacy region of the intersection of the $n$ $(\varepsilon_{i},\delta_{i})$-DP guarantees by
@@
In other words, this intersection region signifies the privacy region of a mechanism with trade-off function $f_{\boldsymbol{\varepsilon},\boldsymbol{\delta}}$. We study the $k$-fold composition of such mechanisms.
@@ II-C Mixture of Hypothesis Tests
Consider that $f_{m}(t)=\inf_{\hat{\mathcal{H}}_{m}}\left\{\beta_{\mathrm{II}}^{m}|\beta_{\mathrm{I}}^{m}\leq t\right\}$. For $\nu\in\left\{\mathrm{I},\mathrm{II}\right\}$, $\beta_{\nu}(\hat{\mathcal{H}}_{m})=\sum_{i=1}^{n}\alpha_{i}\beta_{\nu}^{i}(\hat{\mathcal{H}}_{m}(\cdot,i)):=\sum_{i=1}^{n}\alpha_{i}\beta_{\nu}^{i}(\hat{\mathcal{H}}^{i})$ by routine arguments. The reformulated constraint becomes $\beta_{\mathrm{I}}^{m}=\sum_{i=1}^{n}\alpha_{i}t_{i}\leq t$, $t_{i}=\beta_{\mathrm{I}}^{i}(\hat{\mathcal{H}}^{i})$. Then observe that $f_{i}(t_{i})=f_{i}(\beta_{\mathrm{I}}^{i}(\hat{\mathcal{H}}^{i}))\leq\beta_{\mathrm{II}}^{i}(\hat{\mathcal{H}}^{i})$, leading to the rewriting of $\beta_{\mathrm{II}}^{m}$ as $\sum_{i=1}^{n}\alpha_{i}f_{i}(t_{i})$. The precise arguments can be found in Appendix -A. ∎
@@
Equation (7) is known in the convex optimization literature as a weighted infimal convolution. Its properties are studied in [4], from which the following corollary becomes apparent.
@@
See Figure 1 for an example of the mixture of two equally weighted DP regions.
@@ III-A Heterogeneous Composition
We state the theorem on heterogeneous composition (cf. Definition 4) and the accompanying Algorithm 1.
@@ III-B Composition Theorem for Double-DP Constraints
A first upper bound of the double-DP $k$-composition privacy region is the intersection of the privacy regions for each DP guarantee, i.e. $\mathcal{R}_{k}\left(\boldsymbol{\varepsilon},\boldsymbol{\delta}\right)\subseteq\mathcal{R}_{k}\left(\varepsilon_{1},\delta_{1}\right)\cap\mathcal{R}_{k}\left(\varepsilon_{2},\delta_{2}\right)$, known thanks to [17].
Remark 1 gives a relatively trivial baseline for our main composition result. This baseline can be improved by leveraging the total variation bound induced by DP, as follows:
@@
As discussed in [13]: Let $\eta=\delta_{2}+(1-\delta_{2})\frac{e^{\varepsilon}_{2}-1}{e^{\varepsilon}_{2}+1}$. Then, $\mathcal{R}_{k}\left(\boldsymbol{\varepsilon},\boldsymbol{\delta}\right)\subseteq\mathcal{R}_{k}\left(\tilde{\boldsymbol{\varepsilon}},\tilde{\boldsymbol{\delta}}\right)\cap\mathcal{R}_{k}\left(\varepsilon_{2},\delta_{2}\right)$, where $\tilde{\boldsymbol{\varepsilon}}=(\varepsilon_{1},0)$ and $\tilde{\boldsymbol{\delta}}=(\delta_{1},\eta)$.
In our main composition results, we derive the exact double-DP composition privacy region which improves on Remarks 1 and 2.
@@
$P_{i}(x)=\begin{cases}(1-\alpha)P_{i}^{1}(x)&\text{ if }x\in\left\{0,3\right\},\\ \alpha P_{i}^{2}(x)&\text{ if }x\in\left\{1,2\right\}.\end{cases}$
@@
See Figure 2 for comparison of the bounds given by Remarks 1 and 2 with the exact privacy region computed in Theorems 2 and 3.
@@ IV A Method for Approximating $f$-DP Composition
In the following, we algorithmically approximate the $k$-composition of a trade-off function $f$, yielding both lower and upper privacy bounds***The corresponding code is public on github at https://go.epfl.ch/multiDP.. This is done by approximating $f$ itself from below by a double-DP trade-off function $f_{\boldsymbol{\varepsilon}^{-},\boldsymbol{\delta}^{-}}$, and same from above by $f_{\boldsymbol{\varepsilon}^{+},\boldsymbol{\delta}^{+}}$, then computing the $k$-composition of each using Theorems 2 or 3. The function $f_{\boldsymbol{\varepsilon}^{-},\boldsymbol{\delta}^{-}}$ is defined by
@@ IV-A $f$-DP approximation from below
Below, we define the normal rotation of $f$. It is the rotation of the graph of $f$ by $\frac{\pi}{4}$, changing the axes of the coordinate system from $y=0$ and $x=0$ to y=xy=-x and $y=x$. This allows the removal of the self-symmetry constraint, turning it into an evenness constraint in the rotated domain, which is easier to deal with.
@@ IV-C Example: Gaussian Mechanism
In Figure 3, we obtain empirically close lower and upper bounds on $1$-GDP compositions through double-DP, which will become arbitrarily tight once we generalize our results to $n$-DP composition, as $n$ increases.
@@ V Concluding Remarks
Note that FFT based methods [19, 14] can already achieve arbitrarily close bounds, but are harder to use and more computationally expensive.
Building on the framework introduced in Theorem 2, the generalization to $n$ DP constraints seems natural and will be the subject of future work. With $\boldsymbol{j}=(j_{1},\ldots,j_{n})\in(\mathbb{N}^{*})^{n}$, we refer to the privacy region of the heterogeneous composition of $j_{i}$ $\varepsilon_{i}$-DP for $i\in\llbracket 1\mkern 1.5mu,\mkern 1.5mun\rrbracket$ by $\mathcal{C}_{\boldsymbol{j}}\left(\boldsymbol{\varepsilon}\right)$. Using Corollary 1 and our insight resulting from the case $n=2$ in Theorem 2, we conjecture that we can represent $\mathcal{R}_{k}\left(\boldsymbol{\varepsilon},\boldsymbol{\delta}\right)$ as a convex combination of heterogeneous composition regions $\mathcal{C}_{\boldsymbol{j}}\left(\boldsymbol{\varepsilon}\right)$ and $\mathcal{R}\left(0,1\right)$.